if a right triangle has a hypotenuse of 13cm and one leg that measures 12cm.what is the length of the other leg?

Answer by stanbon(75887)

if a right triangle has a hypotenuse of 13cm and one leg that measures 12cm.what is the length of the other leg?
--------
13^2 = x^2 + 12^2
x^2 = 169-144
x = 5cm
========================
Cheers,
Stan H.